Suppose that the monetary base is $ 500 billion, the reserve deposit ratio is 0.1 and the currency-deposit ratio is 0.6Calculate: i. Money Multiplier (2 Marks) ii. Money Supply


1
Expert's answer
2021-08-12T17:41:41-0400

Part i

Money multiplier "= \frac{1}{ reserve \space deposit \space ratio} = \frac{1}{ 0.1} = 10"

Part ii

Money supply = "Money \space multiplier \times total \space deposit = 10 \times 500\space billion = 5000 \space billion"